OpenStudy (anonymous):

find the product -7(-a^2)^2(-b^3)

OpenStudy (anonymous):

\((-a^2)^2=(-a^2)(-a^2)=a^4\)

OpenStudy (anonymous):

all together therefore you would get \[7a^4b^3\]